LOG-IN
Displaying reports 341-360 of 3360.Go to page Start 14 15 16 17 18 19 20 21 22 End
R&D (Cryogenic)
Print this report.
RishabhBajpai - 09:07, Wednesday 17 April 2024 (3529)Get code to link to this report
Vacuum Pumping

I turned on the scroll pump today for chamber evacuation. I did not notice any issue with the scroll pump operation.

  • Turned on Scroll Pump at 8:03
  • Opened right angle valve between TMP and scroll pump at 8:08
  • Opened gate valve between chamber and TMP at 8:15 (fully open at 8:18)
  • Turned on Vacuum guage at 8:23 (The pressure was 3.75*10Pa)
Comments related to this report
RishabhBajpai - 13:49, Wednesday 17 April 2024 (3533)

I turned on the TMP at 12:54. The pressure at that time was 3.58*101 Pa.

The pressure at 13:05 was 8.1*10-2 Pa.

I will continue pumping for a while (until Monday).

Note: The TMP is in "local" mode and can only  be operated using the front control panel.

RishabhBajpai - 08:29, Monday 22 April 2024 (3535)

The pressure at 8:00 was 4.7*10-4 Pa. Looking at previous logs, the final pressure is consistent with past values.

I plan to start cooling on Thursday.

I also checked the valves for water supply to compressor are open. So I might not have to worry about cleaning up the water lines. I heard the water is supplied by ATC, but do I need to turn on the water chiller?

BIGFOOT (General)
Print this report.
HugoSaintOlive - 09:27, Tuesday 16 April 2024 (3528)Get code to link to this report
Manipulation 2D polarization
After raising the beam with the help of the mirrors at angle, I have characterized the beam obtained with the Beam profiler and a lens f=30mm.
I find z0 = 31cm et w0 = 9.1μm, (z=0 : Corresponds to the characterization lens --> 79,7cm on the meter of the table).
So I found the characteristics of the initial beam : z0 = -9.915m et w0 = 272μm (z=0 still the same).

After that I rebuilt a telescope with two lenses : F= 500mm Bi-Convexe et F=-25,4 Plano-Concave, second one place at 500 mm because the focal seems to be at 525mm...
I put the single pixel Camera at the end, just after the second lens.

Finally, I put the first LCD in front of the large beam just after the mirrors, and I have a control of the monitor, and it seems to work but several remarks :
-There is an important diffraction of the light due to the array of the LCD ---> important power losses
-I remark that we still have color on the LCD, so we don't only have the LC on the screen...
-We need a second computer to put the image on the screen because we have no idea what we put on the screen except when you use two polarizers and a flashlight to see what is on the screen.
R&D (Cryogenic)
Print this report.
RishabhBajpai - 18:56, Monday 15 April 2024 (3527)Get code to link to this report
Closing Chamber

With help from Takahashi-san, I closed the vacuum chamber.

I didn't start pumping since I would like to monitor the operation of the pumps over a few hours from a safety point of view (since they haven't been operated in quite a while). Also, I haven't set up any DAQ system to log the pressure. I will start pumping this Thursday or next Monday.

R&D (Cryogenic)
Print this report.
RishabhBajpai - 15:57, Wednesday 10 April 2024 (3526)Get code to link to this report
Vacuum Fittings

Overview: I modified the vacuum connections around the TMP. 

Before: The TMP is located at the bottom of the cryostat with a gate valve and 6-way cross fitting in between. The cross fitting has CF (-250?) flanges on either side with four "KF ports" on the cylindrical body. The "KF ports" are located after the gate valve. The outlet port of TMP is KF16 and connected to a "KF16-KF25 adaptor". This adaptor is connected to a "Right angle valve", which is connected to a "KF25 TEE" fitting. One end of the "TEE" fitting should be connected to the scroll pump with a flexible vacuum hose. The other end of the "TEE" was attached to one of the "KF ports" with another "Right angle valve". The remaining "KF ports" had a vacuum gauge, a Swagelock poppet check valve and an inlet valve.

Changes: I replaced the swagelock poppet check valve with a blank flange. Furthermore, the "Right angle valve" on one of the "KF ports" connected to the "TEE" with a flexible vacuum hose was removed and replaced with a blank flange. The "Right angle valve" was connected to the "TEE" with a "KF25-Swagelock adaptor" connected to the other end. This adaptor will be used to inject Helium during warm-up.

Items: For this work, I brought the following items from TAMA: 2*KF25 Blank flanges, 3*KF25 hinge clamps, 5*KF25 O-Rings and 1*KF25 full nipple fittings. Additionally, the "KF25-Swagelock adaptor" was brought from KEK.

NOTE: I noticed the KF16 outlet port was not tightened. I tried to tighten it, but it was still loose. Not sure if this is normal.

R&D (Cryogenic)
Print this report.
RishabhBajpai - 08:18, Wednesday 10 April 2024 (3525)Get code to link to this report
Temperature Sensor

I installed a temperature sensor (brought from KEK) on the breadboard inside the cryostat (photo). It is connected to wire bundle C2. I think there is a temperature sensor mounted on the cold-head but I can't find which pin feedthrough it is connected to (this wire perhapas).

KAGRA MIR (Absorption)
Print this report.
MarcEisenmann - 10:14, Saturday 06 April 2024 (3524)Get code to link to this report
Comment to T120#1 (Click here to view original report: 3520)

z = 74.5,mm, Pt = 4.81W : Sat, Apr 6, 2024 10-14-17 AM.txt

KAGRA MIR (Absorption)
Print this report.
MarcEisenmann - 21:53, Friday 05 April 2024 (3523)Get code to link to this report
Comment to T120#1 (Click here to view original report: 3520)

z = 44.5, Pt = 4.87W : Fri, Apr 5, 2024 9-52-57 PM.txt

KAGRA MIR (Absorption)
Print this report.
MarcEisenmann - 17:13, Friday 05 April 2024 (3522)Get code to link to this report
Comment to T120#1 (Click here to view original report: 3520)

z = 54.5mm, Pt = 4.87W ; Fri, Apr 5, 2024 5-12-55 PM.txt

KAGRA MIR (Absorption)
Print this report.
MarcEisenmann - 13:14, Friday 05 April 2024 (3521)Get code to link to this report
Comment to T120#1 (Click here to view original report: 3520)

z edges seem to be at 39 and 90 mm. z_center = 64.5mm

The DC is still slowly decreasing with increasing z from aboiut 5 to 3 V... depending on measurements results I might aligned DC to specific z.

map at z_center : Fri, Apr 5, 2024 1-13-49 PM.txt

KAGRA MIR (Absorption)
Print this report.
MarcEisenmann - 11:53, Friday 05 April 2024 (3520)Get code to link to this report
T120#1

Installed the 30mm diameter , 120mm thickness sample (1623048-4) called 120#1

I moved z_IU to 18.51mm

pink tape is at bottom towards IU

the sample is held roughly t its central thickness position to limit its tilt that could misaligned the beam on the sensors.

in that configuration, the DC is still varying between about 2 to 5V while changing z from 20 to 120mm...

Pt = 27.0mW ; Pin = 31.3mW

Pt = 4.72W

long z scan at x,y centers : Fri, Apr 5, 2024 11-52-43 AM.txt

Comments related to this report
MarcEisenmann - 13:14, Friday 05 April 2024 (3521)

z edges seem to be at 39 and 90 mm. z_center = 64.5mm

The DC is still slowly decreasing with increasing z from aboiut 5 to 3 V... depending on measurements results I might aligned DC to specific z.

map at z_center : Fri, Apr 5, 2024 1-13-49 PM.txt

MarcEisenmann - 17:13, Friday 05 April 2024 (3522)

z = 54.5mm, Pt = 4.87W ; Fri, Apr 5, 2024 5-12-55 PM.txt

MarcEisenmann - 21:53, Friday 05 April 2024 (3523)

z = 44.5, Pt = 4.87W : Fri, Apr 5, 2024 9-52-57 PM.txt

MarcEisenmann - 10:14, Saturday 06 April 2024 (3524)

z = 74.5,mm, Pt = 4.81W : Sat, Apr 6, 2024 10-14-17 AM.txt

R&D (FilterCavity)
Print this report.
MichaelPage - 00:46, Friday 05 April 2024 (3518)Get code to link to this report
OPO replacement PZT

During Yuhang's Christmas visit we attempted to replace the TAMA OPO. It was aligned but could not take a scan signal - the high voltage driver would just zero out then turn off, suggesting that the PZT on the input mirror is broken. I have to disassemble the "new" OPO and replace the PZT. The connectors for the PZT are soldered to our Dsub box - the OPO was kept in a zip lock bag during soldering. The Dsub box connects the OPO PZT and temperature actuator wires and puts them on a Dsub port, which then plugs into the controls of TAMA electronics.

I measured the output power vs current of the Coherent Mephisto laser in the ATC cleanroom used for the Taiwan and Korea experiments, using both standard Thorlabs (up to 500 mW) and thermal (several W) power meters. I fixed the laser temperature at 24.21 C. The output of the laser is about 47 mW at 0.8 A, up to 1.17 W at 1.5 A (it can go higher but I stopped here).

In the leadup to the Taiwan/Korean Christmas visit I designed the input beam path to: 1) inject collimated infrared p-polarized light into the Taiwan fiber SHG (~2.3mm diameter), and 2) send a small amount of s-polarized light to the Korea Faraday isolator. The actual beam profile was measured by Yi Ru Chen. There is a HWP and PBS setup which transmits p-pol to Taiwan (SHG) and reflects s-pol to Korea (Faraday). I set up to reflect a small amount of s-pol to Korea. At 0.800 A laser current, 30 mW reaches the HWP, with 3.8 mW s-pol going to Korea and 25.1 mW p-pol going to Taiwan. 3.8 mW is reasonable for our OPO characterization, it's about what we used last time. There is plenty of room to adjust up or down anyway, and since the beam size going to the OPO is around 20 micron the intensity is quite dangerous even at 4 mW.

I constructed a beam path to travel across the indicated line on the breadboard, cutting off from the Korean FI path (they aren't going to come back until October). The beam travels horizontally at 75mm, the height of the OPO crystal. In a previous iteration of this test we had a rotation stage below the OPO for rotation alignment but we found it to be unnecessary. It is better to set a beam at constant height, rough align the OPO case by hand then fine align with steering mirrors. We also did it this way in TAMA twice during the replacement/re-replacement episode and it works fine, could achieve > 90 % mode matching quickly. The beam path should be collimated at 2.3mm for quite some distance, so based on this I can set up a system to send the beam through an EOM. The lenses we originally used to mode match to the OPO (40 and 75 mm) are in the ATC unused.

I took a soldering iron from the elec shop and the Beam Master beam profiler from birefringence compensation. For other instruments the ATC KOACH cleanroom now has the 8 GHz function generator, PZT high voltage driver, Thorlabs temperature controller, resonant EOM, oscilloscope, old TAMA spectrum analyzer, PDA05CF2 RFPD, DET10N/M biased PD. We can scan the PZT of the reassembled OPO to test mode matching but to lock I would need another function generator or mokulab. The CCD and TV screen were taken to speed meter but maybe I don't need them. Locking is probably unnecessary though. It seems based on previous calculations that, due to the short length and low finesse, the storage time is too short to determine internal loss from cavity ring-down, even with a nanosecond scale off-switch. We will need to wait until the sqz/antisqz measurement to really see the optical loss.

KAGRA MIR (Absorption)
Print this report.
MarcEisenmann - 22:28, Thursday 04 April 2024 (3519)Get code to link to this report
Comment to T70#3 (Click here to view original report: 3514)

z = 73.75mm, Pt = 4.87W : Thu, Apr 4, 2024 10-28-30 PM.txt

KAGRA MIR (Absorption)
Print this report.
MarcEisenmann - 18:31, Thursday 04 April 2024 (3517)Get code to link to this report
Comment to T70#3 (Click here to view original report: 3514)

z = 53.75mm, Pt = 4.85W : Thu, Apr 4, 2024 6-30-57 PM.txt

KAGRA MIR (Absorption)
Print this report.
MarcEisenmann - 14:24, Thursday 04 April 2024 (3515)Get code to link to this report
Comment to T70#3 (Click here to view original report: 3514)

edges at 48.5 and 79 mm : z_center = 63.75 mm

map at z_center : Thu, Apr 4, 2024 2-23-48 PM.txt

KAGRA MIR (Absorption)
Print this report.
MarcEisenmann - 11:59, Thursday 04 April 2024 (3514)Get code to link to this report
T70#3

I installed the 30mm diameter; 70mm length, 4/6 sapphire (T70#3) with pink tape at bottom towards IU.

Realigned DC

Pt = 26.3mW ; Pin = 30.7 mW

Pt = 4.87W ; long z scan (as for all previous z scan at x,y center ie 327,119 mm) : Thu, Apr 4, 2024 11-59-20 AM.txt

Comments related to this report
MarcEisenmann - 14:24, Thursday 04 April 2024 (3515)

edges at 48.5 and 79 mm : z_center = 63.75 mm

map at z_center : Thu, Apr 4, 2024 2-23-48 PM.txt

MarcEisenmann - 18:31, Thursday 04 April 2024 (3517)

z = 53.75mm, Pt = 4.85W : Thu, Apr 4, 2024 6-30-57 PM.txt

MarcEisenmann - 22:28, Thursday 04 April 2024 (3519)

z = 73.75mm, Pt = 4.87W : Thu, Apr 4, 2024 10-28-30 PM.txt

KAGRA MIR (Absorption)
Print this report.
MarcEisenmann - 22:00, Wednesday 03 April 2024 (3513)Get code to link to this report
Comment to t70#2 (Click here to view original report: 3509)

z = 66.5mm, Pt = 4.83W : Wed, Apr 3, 2024 9-59-53 PM.txt

R&D (FilterCavity)
Print this report.
MarcEisenmann - 20:19, Wednesday 03 April 2024 (3512)Get code to link to this report
Comment to tidying up electronics (Click here to view original report: 3508)

I swapped the R and L cables for the OPO mixer.

During this process a cable for CC PLL monitor was quite damagedd but it's a bit tedious to replace alone and not priority for now so I left it as is for now.

I finalized the new box connections and little cable tidy up.

I could lock SHG after changing the DDS1 DAC1 phase from 20 to 240deg and the servo gain from 0.9 to 0.1 to avoid ringing. I measured about 30% HOM still present and 235mW of green produced.

The MZ/GRMC are heavily misaligned and I'll recover them tomorrow.

BIGFOOT (General)
Print this report.
HugoSaintOlive - 19:19, Wednesday 03 April 2024 (3511)Get code to link to this report
Manipulation 2D polarization
I choose and align two different lenses in order to extend my initial beam : - 1 inch ; f=25.4mm . - 2 inches ; f=750mm.

I've installed my two supports, made using a 3D printer, for the LCD (from the screen of a monitor).

The beam of the setup is too low to path throw the LCD screen. So I added to more mirrors (2 inches) to get the beam higher, and I started to align them.
KAGRA MIR (Absorption)
Print this report.
MarcEisenmann - 17:28, Wednesday 03 April 2024 (3510)Get code to link to this report
Comment to t70#2 (Click here to view original report: 3509)

z = 46.5m, Pt = 4.73W : Wed, Apr 3, 2024 5-28-27 PM.txt

KAGRA MIR (Absorption)
Print this report.
MarcEisenmann - 10:32, Wednesday 03 April 2024 (3509)Get code to link to this report
t70#2

Installed T70#2 (30; 70 mm 5/6)

pink tape at bottom towardsiu

realigned DC

Pt = 27.4mW ; Pin = 31.5mW

long z scan Pt = 4.77W ; Wed, Apr 3, 2024 12-31-34 PM.txt

z edges at 36 and 77 mm : z_center = 56.5 mm

map at z_center : Wed, Apr 3, 2024 1-11-48 PM.txt

Comments related to this report
MarcEisenmann - 17:28, Wednesday 03 April 2024 (3510)

z = 46.5m, Pt = 4.73W : Wed, Apr 3, 2024 5-28-27 PM.txt

MarcEisenmann - 22:00, Wednesday 03 April 2024 (3513)

z = 66.5mm, Pt = 4.83W : Wed, Apr 3, 2024 9-59-53 PM.txt